A mother is twice as old as her daughter now, but 10 years ago, she was 3 times as
old as her daughter. Find their ages at present.

Answer by mananth(12270)

Let the daughter's age be x
mother's age will be 2x
Ten years ago
daughter's age x-10
Mother's age 3(x-10)
x-10 =3(x=10)
x=10=3x-30
3x-x=10+30
2x=40
x=40/2
x=20 Daughter's age now
Mother's age is 2x = 2*20 = 40 years